# Towards Assessing Medical Ethics from Knowledge to Practice

### v1

@@ Abstract
The integration of large language models into healthcare necessitates a rigorous evaluation of their ethical reasoning, an area current benchmarks often overlook. We introduce PrinciplismQA, a comprehensive benchmark with 3,648 questions designed to systematically assess LLMs’ alignment with core medical ethics. Grounded in Principlism, our benchmark features a high-quality dataset. This includes multiple-choice questions curated from authoritative textbooks and open-ended questions sourced from authoritative medical ethics case study literature, all validated by medical experts. Our experiments reveal a significant gap between models’ ethical knowledge and their practical application, especially in dynamically applying ethical principles to real-world scenarios. Most LLMs struggle with dilemmas concerning Beneficence, often over-emphasizing other principles. Frontier closed-source models, driven by strong general capabilities, currently lead the benchmark. Notably, medical domain fine-tuning can enhance models’ overall ethical competence, but further progress requires better alignment with medical ethical knowledge. PrinciplismQA offers a scalable framework to diagnose these specific ethical weaknesses, paving the way for more balanced and responsible medical AI.
@@ 1 Introduction
Large language models (LLMs) have emerged as a transformative force in artificial intelligence, generating significant interest across diverse domains, including healthcare (Singhal et al., 2023; Guha et al., 2023; M. Bran et al., 2024). In the medical domain, LLMs show promise for applications such as clinical decision support (Singhal et al., 2025; Goh et al., 2024) and patient communication (Ayers et al., 2023; Liu et al., 2025). However, despite encouraging early results, the field remains in its infancy (Thirunavukarasu et al., 2023; Clusmann et al., 2023). Given the complexity of medical knowledge and the critical importance of patient safety, rigorous and systematic evaluation of these models is imperative. To ensure their responsible integration into clinical practice, standardized assessment frameworks must be established to balance potential benefits against risks. Thus, advancing robust evaluation methodologies demands urgent attention.
Medical LLM evaluations often prioritize diagnostic accuracy and knowledge retrieval, neglecting crucial aspects like adherence to structured clinical workflows and medical ethics. Real-world decision-making relies on deterministic constraints (e.g., FDA guidelines, triage protocols) and ethical principles, including patient autonomy, informed consent, equity in care, and prioritization in resource-limited settings. Current frameworks rarely address these ethical dimensions, leaving gaps in evaluating models’ real-world applicability and ethical soundness. Integrating LLMs into medical ethics is a novel concept, and understanding the effectiveness of these models in aiding ethicists with decision-making can have significant implications for the healthcare sector.
Motivated by these challenges, we designed PrinciplismQA, a medical ethics benchmark, with two primary evaluation objectives:
  1. 1.
    Knowledge Readiness: Through multiple-choice questions (MCQA), we assess whether the model possesses relevant medical ethical knowledge and evaluate its understanding and recall of established ethical principles and guidelines.
  2. 2.
    Human Value Alignment: We use open-ended questions based on real-world clinical ethical dilemmas. For each case, a checklist of essential ethical reasoning points is constructed from peer-reviewed expert commentary, representing consensus human values in medical ethics. By evaluating whether LLM responses address these checklist items, we assess how well the model’s reasoning aligns with human ethical standards. This is automated via the LLM-as-a-Judge paradigm.
The key contributions of this work are as follows.
  • Principlism-based Evaluation Protocol: We leverage Principlism, a widely recognized theoretical framework in medical ethics, to design an evaluation protocol that systematically analyzes both the ethical knowledge and value alignment of LLMs across overall and individual Principlism dimensions.
  • Simulated Clinical Examination Process: PrinciplismQA simulates the clinical evaluation process by incorporating checklist-based methods from real-world medical examinations, enabling innovative and automated assessment of LLMs’ practical ethical reasoning.
  • A Domain-Specific Medical Ethics Dataset: We construct a high-quality, evidence-based dataset explicitly tailored for the evaluation of medical ethics in LLMs, with each question independently reviewed and validated by human medical ethics experts. This provides a robust foundation for benchmarking and further research in this critical area.
2 Philosophy
2.1 Principlism on Medical Ethics
Ethics is an inherent and inseparable part of clinical medicine (Singer et al., 2001) as the physician has an ethical obligation (i) to benefit the patient, (ii) to avoid or minimize harm, and to (iii) respect the values and preferences of the patient. In 1979, Tom Beauchamp and James Childress popularized the use of Principlism in efforts to resolve ethical issues in clinical medicine (Beauchamp and Childress, 2019),
  • Autonomy: Respecting a patient’s right to make their own informed decisions about their healthcare, including the right to refuse treatment.
  • Non-Maleficence: "Do no harm." Avoiding actions or treatments that may cause unnecessary harm or suffering to a patient.
  • Beneficence: Acting in the best interest of the patient by providing care that maximizes benefits and promotes well-being.
  • Justice: Ensuring fair distribution of healthcare resources, equal treatment for all patients, and ethical decision-making in allocation and access to medical services.
In March 2025, the World Health Organization (WHO) released guidelines for the ethical governance of LLMs in clinical care (Wang et al., 2024). While these guidelines currently treat LLMs as auxiliary tools, they anticipate a future where LLMs may act as independent diagnostic agents. This vision highlights the urgent need for rigorous, domain-specific ethical evaluation.
@@ 2.1 Principlism on Medical Ethics
Mitigating Risks Are the risks of harm, such as "hallucinations" (incorrect or misleading information) or biases, effectively mitigated?
@@
Equitable Access Does the LLM pass when providing educational content related to the ethical principle of "Justice"?
@@
During benchmarking, we systematically reviewed how explicit knowledge from medical ethics textbooks relates to real-world cases in medical journals (Ong et al., 2024b). The ultimate expectation for LLMs is to function as independent diagnosticians, demonstrating decision-making abilities comparable to those of clinical medicine students. Accordingly, Table 1 was carefully designed to map each PrinciplismQA question to its corresponding scenario and to the type of scenario in which the LLM would apply its knowledge.
2.2 Principlism Evaluation with Knowledge and Practice Medical Exams
The evaluation framework of PrinciplismQA integrates Principlism as a guiding structure, analogous to the dual approach of the Objective Structured Clinical Examination (OSCE), a gold standard in medical competency assessment that combines both theoretical knowledge and practical skills (Zayyan, 2011). Reflecting this, PrinciplismQA comprises two complementary categories, Knowledge and Practice, the main differencies of them are listed in Table 2.
Table 2: PrinciplismQA Knowledge vs. Practice: Key Characteristics
Knowledge
Purpose: Assess foundational ethical knowledge
Format: MCQAs
Focus: Principlism definitions, concepts, guidelines
Metrics: Accuracy
Assessment: Knowledge recall, understanding
Analogy: Theory test
Practice
Purpose: Assess practical reasoning and decision-making
Format: Open-ended cases
Focus: Application of principles to scenarios
Metrics: Coherence, depth, principle alignment
Assessment: Reasoning, judgment, context application
Analogy: Simulated scenario
Category I: Knowledge evaluates an LLM’s grasp of the foundational concepts and established guidelines underpinning Principlism through multiple-choice questions (MCQAs). This part focuses on measuring the model’s ability to recall and understand ethical definitions, core principles, and consensus views in the medical context.
Category II: Practice assesses the LLM’s reasoning and decision-making in nuanced, real-world clinical ethics scenarios. Here, open-ended cases are used to determine how effectively the LLM can apply principlist concepts to practical dilemmas, simulating the type of complex judgment required in actual medical settings.
Existing medical ethics evaluations focus on knowledge but overlook practice. Recent works map the ethical challenges of LLMs in medicine, focusing on transparency, bias, fairness, and clinician or stakeholder perspectives (Haltaufderheide and Ranisch, 2024; Gallegos et al., 2024; Mirzaei et al., 2024; Ong et al., 2024a; Pressman et al., 2024). Some probe the limits of LLMs in ethical vignettes (Balas et al., 2024), but do not offer systematic, principle-based benchmarking. Existing benchmarks such as MedSafetyBench (Han et al., 2024) emphasize safety refusal based on AMA standards, overlooking nuanced ethical reasoning in real-world clinical cases. MedEthicEval (Jin et al., 2025) is narrowly scoped, limited to self-defined scenarios, and primarily assesses Chinese-language responses, lacking multilingual and international breadth. Other datasets, such as MedEthicsQA (Wei et al., 2025), recognize the importance of Principlism but do not conduct sufficient open-ended practice evaluations that capture practical complexity. Large-scale resources (Bian et al., 2025) also tend to focus on governance frameworks over principled ethical reasoning. Collectively, these gaps highlight the lack of a comprehensive, internationally relevant benchmark that robustly evaluates LLMs’ ethical reasoning across diverse clinical scenarios and all principlism-based ethical principles, rather than just safety or recall.
3 PrinciplismQA
PrinciplismQA is designed to evaluate Principlism biases in medical ethics reasoning of large language models. It comprises two task-oriented subsets Knowledge and Practice, where the former contains 2182 knowledge MCQAs, while the latter involves 1466 open-ended questions. Knowledge questions were derived from 350 authoritative medical ethics textbooks in English published globally since 2010. Practice includes real-world clinical ethical cases sourced from JAMA’s Journal of Ethics, each with structured ethical checklists based on expert peer-reviewed commentaries. Questions are labeled across four ethical principles from Principlism—autonomy, beneficence, non-maleficence, justice—and validated by medical experts, ensuring dataset accuracy, diversity, and clinical relevance.
Table 3: Question Distribution by Ethical Principle
@@ 3 PrinciplismQA
Notes. “Multiple principles*” indicates questions involving more than one principle.
Table 3 presents both multiple-choice (MCQA) and open-ended questions categorized by the four fundamental ethical principles. Notably, 285 MCQAs (13.1%) involved multiple ethical concerns, reflecting the complexity of real-world medical ethics reasoning. Moreover, 852 open-ended questions(58.1%) involved multiple ethical principles, highlighting the complex ethical dilemmas frequently encountered in medical practice.
3.1 Data Curation and Annotation
Figure 1 presents the complete PrinciplismQA benchmarking workflow, comprising two phases: data curation and annotation.

### v2

# PrinciplismQA: A Philosophy-Grounded Approach to Assessing LLM-Human Clinical Medical Ethics Alignment

@@ Abstract
As medical LLMs transition to clinical deployment, assessing their ethical reasoning capability becomes critical. While achieving high accuracy on knowledge benchmarks, LLMs lack validated assessment for navigating ethical trade-offs in clinical decision-making where multiple valid solutions exist. Existing benchmarks lack systematic approaches to incorporate recognized philosophical frameworks and expert validation for ethical reasoning assessment. We introduce PrinciplismQA, a philosophy-grounded approach to assessing LLM clinical medical ethics alignment. Grounded in Principlism, our approach provides a systematic methodology for incorporating clinical ethics philosophy into LLM assessment design. PrinciplismQA comprises 3,648 expert-validated questions spanning knowledge assessment and clinical reasoning. Our expert-calibrated pipeline enables reproducible evaluation and models ethical biases. Evaluating recent models reveals significant ethical reasoning gaps despite high knowledge accuracy, demonstrating that knowledge-oriented training does not ensure clinical ethical alignment. PrinciplismQA provides a validated tool for assessing clinical AI deployment readiness.
PrinciplismQA: A Philosophy-Grounded Approach to Assessing LLM-Human Clinical Medical Ethics Alignment
Chang HONG1⋆, Minghao WU1⋆, Qingying XIAO2‡, Yuchi WANG1, Xiang WAN3‡, Guangjun YU1,2, Benyou WANG1, Yan HU2 1The Chinese University of Hong Kong, Shenzhen 2National Health Data Institute, Shenzhen  3Shenzhen Research Institute of Big Data {changhong, minghaowu, yuchiwang}@link.cuhk.edu.cn {xiaoqingying, wanxiang, guangjunyu, wangbenyou, huyan}@cuhk.edu.cn Equal Contribution   Corresponding Author
@@ 1 Introduction
Medical LLMs now achieve high accuracy on benchmarks such as USMLE-like MedQA Jin et al. (2021) and open-ended question-focused HealthBench Arora et al. (2025), which focus on identifying “one of the valid solutions”. This high performance demonstrates apparent deployment readiness. However, ground truth-oriented benchmark paradigms create a paradox between technological capability and ethical considerations.
Current ethical assessments of LLMs concentrate on AI safety Gallegos et al. (2024); Ong et al. (2024) mechanisms such as privacy protection and automatic personally identifiable information (PII) data masking. Unlike these well-defined safety tasks, practical clinical dilemmas involve navigating conflicting ethical principles across multiple valid solutions, which we term “multiple-to-one” decision-making (see Figure 1). Most LLMs, including state-of-the-art (SOTA) models, typically propose a single solution and demonstrate its validity rather than explicitly comparing alternatives. Medical ethics considerations remain largely absent from their selection process.
Figure 1: Distinguishing “valid solution” identification from ethical deliberation. Human clinicians explicitly compare alternatives using ethical principles, while LLMs default to frequent training patterns without comparative analysis, revealing a critical gap between benchmark performance and deployment readiness.
@@
Mitigating Risks Are the risks of harm, such as “hallucination” (incorrect or misleading information) or biases, effectively mitigated?
@@
Equitable Access Does the LLM pass when providing educational content related to the ethical principle of “Justice”?
@@
The assessment paradigm-oriented nature of LLM development reveals three key limitations triggering this absence of medical ethics considerations. First, current medical benchmarks prioritize knowledge recall and clinical reasoning that improve response precision, treating this as a root metric of medical AI. Second, few benchmarks model medical ethics using gold standards, despite its alignment with evidence-based clinical medicine. Third, medical ethics reflects human preferences, requiring medical experts to calibrate benchmarks through clear data protocols and perform secondary verification of evaluation results.
LLMs are increasingly embedded in clinical workflows as documentation assistants, patient communication drafts, and decision-support tools, roles where they already encounter ethically complex scenarios. In these support roles, ethical reasoning gaps become directly consequential. As our Practice subset cases illustrate, most models respond to clinical dilemmas with technically correct information but fail to surface the underlying principle conflicts that clinical ethics consultation would flag. While no model responds perfectly, better-aligned SOTA LLMs do explicitly surface such conflicts, demonstrating that principlist reasoning is measurable and improvable. To increase awareness of incorporating ethical considerations into LLM clinical decision-making and provide a validated tool for bridging this gap, we constructed PrinciplismQA from recognized textbooks and peer-reviewed clinical cases grounded in Principlism Childress and Beauchamp (1994), the gold standard framework in international medical ethics. Through expert validation, we developed a benchmark of 3.6k questions alongside a corresponding assessment pipeline verified for consistency with medical experts.
Benchmark Principlism Complexity Evaluator Scope
MedQA Jin et al. (2021) Diagnosis and Treatment
HealthBench Arora et al. (2025) Clinical reasoning
MedSafetyBench Han et al. (2024) Safety refusal
MedEthicEval Jin et al. (2025) Chinese-language
MedEthicsQA Wei et al. (2025) Knowledge recall
Ethics and Safety QA Bian et al. (2025) Governance
PrinciplismQA (Ours) Clinical deliberation
Table 2: Comparison of medical benchmarks and medical ethics benchmarks. Principlism: whether the benchmark explicitly involves Principlism as assessment philosophy. Complexity: whether the benchmark includes “multiple-to-one” clinical scenarios requiring deep ethical reasoning beyond single solutions or superficial concepts. Evaluator: whether the benchmark provides an evaluation toolkit.
The key contributions of this work are as follows. (1) Philosophy-grounded calibration and validation. We establish procedures and protocols grounded in Principlism, ensuring consistency with established frameworks in clinical practice. This enables systematic assessment against recognized gold standards and supports ethical preference analysis towards each principles from Principlism. (2) Complex clinical scenarios involvement. We introduce scenarios requiring explicit ethical deliberation among multiple valid alternatives, reflecting real-world complexity where clinicians must weigh competing principles to determine optimal care. (3) Expert-validated assessment pipeline. We develop a reproducible evaluation framework validated by medical experts to assess whether LLMs engage in medical ethics considerations when faced with clinical dilemmas.
Through PrinciplismQA and its associated assessment framework, we provide the research community with an approach to measure and improve ethical alignment in medical AI systems, bridging the critical gap between assessment performance and responsible clinical deployment.
2 Philosophy
2.1 Principlism in Clinical Medical Ethics
Ethics is an integral to clinical medicine Singer et al. (2001), as physicians have ethical obligations to benefit patients, avoid or minimize harm, and respect patient values and preferences. In 1979, Tom Beauchamp and James Childress popularized Principlism to resolve clinical ethical issues Beauchamp and Childress (2019), establishing four fundamental principles: (1) Autonomy. Respecting a patient’s right to make informed decisions about their healthcare, including the right to refuse treatment. (2) Non-Maleficence. Avoiding actions or treatments that may cause unnecessary harm or suffering to a patient. (3) Beneficence. Acting in the patient’s best interest by providing care that maximizes benefits and promotes well-being. (4) Justice. Ensuring fair distribution of healthcare resources, equal treatment for all patients, and equitable access to medical services.
Building upon this framework, all protocols for PrinciplismQA are grounded in Principlism. From curation to analysis, each component evaluates whether LLMs navigate Principlism in clinical decisions. This ensures alignment with clinical gold standards during expert verification and provides philosophy-grounded assessment of LLM medical ethics performance.
2.2 Medical Ethics Benchmarks
Recent works have mapped ethical challenges of LLMs in medicine, focusing on transparency, bias, fairness, and stakeholder perspectives Haltaufderheide and Ranisch (2024); Gallegos et al. (2024); Mirzaei et al. (2024); Ong et al. (2024); Pressman et al. (2024). Table 2 summarizes representative medical and medical ethics benchmarks alongside their key characteristics. Among these, two works are closely related yet complementary to PrinciplismQA. MedSafetyBench (Han et al., 2024) evaluates whether LLMs can identify unsafe advice or appropriately refuse malicious queries; by contrast, PrinciplismQA addresses a fundamentally different challenge: principled deliberation among multiple clinically valid options where no single “unsafe” trigger exists. MedEthicsQA (Wei et al., 2025) valuably assesses normative ethical knowledge in abstracted ethical conflicts; PrinciplismQA extends this foundation into real-world clinical dilemmas, where intricate patient histories, conflicting stakeholder demands, and significant ambiguity naturally establish a higher difficulty ceiling.
Figure 2: PrinciplismQA sample with expert-validated assessment rubric. Both discharge and restraint are clinically valid. The rubric evaluates ethical reasoning quality: identification of principle conflicts (autonomy vs. beneficence/non-maleficence), explicit comparison of alternatives, and alignment with expert consensus. Scores reflect comprehensiveness of ethical deliberation rather than binary correctness.
2.3 Research Gaps
As shown in Table 2, three key limitations emerge from current evaluation paradigms.
Lack of philosophy-grounded assessment. Existing medical benchmarks primarily evaluate clinical knowledge and reasoning without systematic grounding in established ethical frameworks. While some benchmarks acknowledge ethical considerations, they lack explicit integration of gold standard frameworks such as Principlism. We address this by grounding PrinciplismQA in the four principles of autonomy, non-maleficence, beneficence, and justice, ensuring alignment with international clinical ethics standards.
Insufficient modeling of clinical complexity. Current benchmarks treat single solutions as correct answers without requiring deep medical ethics considerations. The complexity inherent in clinical ethics, where multiple valid alternatives may exist with different ethical implications, remains largely unmodeled. We address this through clinical cases requiring deliberation among multiple valid solutions, assessing LLM responses based on their explicit consideration of each principle rather than merely identifying a valid option.
Limited reproducibility and validation. Recent benchmarks are published with evaluation toolkits to ensure ease of use and reproducibility. However, the effectiveness of these assessment approaches in capturing nuanced ethical reasoning often lacks expert validation. We follow this trend by developing a corresponding pipeline for PrinciplismQA and validating its assessment effectiveness through medical expert review, ensuring that automated evaluations align with expert consensus on ethical deliberation quality.
3 Constructing PrinciplismQA
3.1 PrinciplismQA Components
PrinciplismQA consists of three integrated components designed to systematically assess LLM ethical reasoning in clinical contexts. First, our philosophy-grounded data engineering protocol provides a systematic methodology for organizing clinical content using the Principlism framework, ensuring all questions are anchored in recognized medical ethics philosophy.
Following this protocol, we curated our benchmark comprising 3,648 questions across two assessment formats: (1) Knowledge questions (2,182 MCQA) that evaluate whether LLMs understand principlist concepts and terminology—serving as the entry criterion for ethical reasoning capability, and (2) Practice questions (1,466 open-ended) that assess whether LLMs can apply principlist reasoning in “multiple-to-one” clinical dilemmas requiring explicit trade-off navigation. As shown in Table 3, practice questions involve substantially higher ethical complexity, with 58.1% requiring navigation of multiple principles simultaneously, compared to 13.1% in knowledge questions. Third, calibrated according to the same protocol, our assessment pipeline (Evaluator), which is a zero-shot agent framework consists of zero-shot candidate LLM module and a SOTA LLM-as-a-Judge scoring module, enables reproducible evaluation through direct answer matching for MCQA and expert-calibrated rubric-based scoring for open-ended questions, addressing the expert validation challenge at scale.
@@ 3.1 PrinciplismQA Components
Table 3: Question distribution according to Principlism. “Multiple principles*” indicates questions involving more than one principle.